OpenStudy (anonymous):

A square has sides of length 5. What is the length of the square's diagonal?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

a^2 +b^2 = c^2 where c is the diagonal and a = 5 and b = 5

OpenStudy (anonymous):

Pythagoras' theorem

OpenStudy (mathstudent55):

Use Pythagoras theorem, a^2 + b^2 = c^2 Here, a = b = 5, c = d d^2 = 5^2 + 5^2

OpenStudy (anonymous):

Thanks :)

OpenStudy (mathstudent55):

wlcm